OpenStudy (anonymous):

A computer is worth $4000 when it is new. After each year it is worth half what it was the previous year. What will its worth be after 5 years? Round your answer to the nearest dollar. Options: $250 $63 $31 OR $125

OpenStudy (anonymous):

It worths $400 after fifth yr.: \[=>\frac{ 4000 }{ 5 \times 2 } = $400\]

OpenStudy (anonymous):

I got 125

OpenStudy (anonymous):

4000 in the first year, divided by two you get 2000 Then 2000/2=100 1000/2=500 500/2=250 250/2=125 that's five years

OpenStudy (anonymous):

Wow, that's a ripofff

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@avalonmay: I also thought the same ...but iin the haste I preffered the unitary method...and I forgot to mention it hetre :: +up